OpenStudy (anonymous):

(3 pt) Cards numbered 1 through 30 are placed in a bag and shuffled. One card is chosen at random. What is the probability of randomly selecting a number less than or equal to 5, or a multiple of 7? A.1/30 B. 2/15 C.1/15 D.3/10

jimthompson5910 (jim_thompson5910):

Make a list of numbers from 1 to 30 Then circle 5 and any multiple of 7 (7, 14, etc) how many circled numbers are there?

OpenStudy (anonymous):

D. \[\frac{5+4}{30}=\frac{3}{10} \]

jimthompson5910 (jim_thompson5910):

Sorry I meant to say "circle numbers less than or equal to 5" instead of just "equal 5"

OpenStudy (anonymous):

@jim_thompson5910 Your doing better than I. Made some big time errors lately. My age, 81 years, is starting to be a real drag.